AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 567     Entry time: Fri May 31 21:03:21 2024
Author: CB, EM 
Subject: Run14 

22:03

Before this run we checked quickly  beam with the target on but it does not make sense with the current status of the beam.

We removed target and starte Run14

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

Attachment 1: Screenshot_from_2024-05-31_22-05-57.png  874 kB  Uploaded Fri May 31 22:06:01 2024  | Hide | Hide all
Screenshot_from_2024-05-31_22-05-57.png
Attachment 2: Screenshot_from_2024-05-31_22-07-01.png  923 kB  Uploaded Fri May 31 22:07:14 2024  | Hide | Hide all
Screenshot_from_2024-05-31_22-07-01.png
Attachment 3: Screenshot_from_2024-05-31_22-07-49.png  602 kB  Uploaded Fri May 31 22:07:55 2024  | Hide | Hide all
Screenshot_from_2024-05-31_22-07-49.png
ELOG V3.1.4-unknown